Dear Financial Adviser, My spouse and I are each 62 and hope to retire in three years. After retirement we will receive $8,500 per month after taxes from our employers pension plans and $2,500 per month after taxes from Social Security. Unfortunately our monthly living expenses are $16,000. Our social obligations preclude further economies. We have $1,100,000 invested in a high-grade, tax-free municipal-bond mutual fund. The return on the fund is 3.0% per year. We plan to make annual withdrawals from the mutual fund to cover the difference between our pension and Social Security income and our living expenses. How many years before we run out of money? Sincerely Luxury Challenged Marblehead, MA You can assume that the withdrawals (one per year) will sit in a checking account (no interest) until spent. The couple will use the account to cover the monthly shortfalls How many years before Luxury Challenged runs out of money? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Number of years

I found 31.08 or 32 years and it says it's wrong.
Answer #1

Amount of Shortfall per month = 16000-(8500+2500) = 5000
Amount of Shortfall per year = 5000*12 = 60000

Amount in Mutual fund = 1,100,000
PMT= 60000
Rate = 3%
Using Excel function for Number of years = NPER(3%,60000,-1100000) = 31.08

However , this takes place after 3 years hence 31.08 + 3 = 34.08 years

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well